OpenStudy (anonymous):

Which of the following shows the length of the third side, in inches, of the triangle below? A right triangle is shown. One side of the triangle is labeled as 61 inches. The height of the triangle is labeled as 11 inches. Square root of 3842 inches 60 inches Square root of 50 inches 72 inches

OpenStudy (welshfella):

Where is the drawing?

OpenStudy (anonymous):

i was thinking B 60

OpenStudy (welshfella):

I can't hep you without the drawing as I don't know f 61 is the longest side or not

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (welshfella):

Right well you use the Pythagoras theorem X^2 = 61^2 - 11^2

OpenStudy (welshfella):

X^2 = 3600

OpenStudy (anonymous):

oh so A?

OpenStudy (welshfella):

Square root of 3600 = 60 Yes B

OpenStudy (anonymous):

yay thank you

OpenStudy (welshfella):

Hold on I haven got a calc with me 61* 61 = 3660+61 = 3721 -121 = 3660 Ye it is B

OpenStudy (welshfella):

Yw

OpenStudy (anonymous):

im back sorry

OpenStudy (anonymous):

oh ok nvm